# Large Language Models for Automated Literature Review: An Evaluation of Reference Generation, Abstract Writing, and Review Composition

### v4

@@ 3.3 Evaluation Metrics
$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}\vspace{-2mm}$ (1)
@@
$Precision_{(p_{i})}=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})\vspace{-2mm}$ (2)
$Preicison=\frac{1}{M}\sum_{i=0}^{M}Precison{(p_{i})}\vspace{-1mm}$ (3)
@@
$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}^{(n)}\vspace{-2mm}$ (4)
$s_{p_{i}}^{(n)}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from external Scholar API,}\\ 0&\text{otherwise}\end{cases}\vspace{-1mm}$ (5)
@@
$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}\vspace{-2mm}$ (7)
@@
$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}\vspace{-2mm}$ (8)
@@ 4.2 Main Results
Table 1: The experimental results of the five LLMs in Reference Generation. “$S_{t}$” refers to the title search rate as defined in Eq 4, while “P” represents the Precision, “R” denotes the Recall. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
@@
Table 3: The experimental results of the four LLMs in Review Composition. “$S_{t}$” refers to the title retrieval rate as defined in Eq 4. While “P” represents the Precision, “R” denotes the Recall. “KPR” means the Key Point Recall rate.“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
@@ Appendix A Prompts for Tasks
Prompt 1 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to find the 10 studies that are most relevant to the research topic provided in the "Title" and the "Keywords" below.
Please cite the studies according to the following JSON format. There is no need to provide any explanation before or after the JSON output. Ensure that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.". { "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
@@
Prompt 2 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write an abstract according to the research topic provided in the "Title" and the "Keywords" below. Please write the abstract for about xx words, according to the JSON format as follows. There is no need to provide any explanation before or after the JSON output. {"Abstract": ""}
@@
Prompt 3 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write a literature review according to the research topic provided in the "Title", “Abstract” and "Keywords" below. The literature review should be about 1000 words long. I would like you to back up claims by citing previous studies (with a total of 10 citations in the literature review). The output should be in JSON format as follows: { "Literature Review": "xxx", "References": [ { "title": "", "authors": "", "journal": "", "year": "", "volumes": "", "first page": "", "last page": "", } ] }
The "Literature Review" field should be about 1000 words. The "References" field is a list of 10 references, and ensures that the "authors" field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as "et al.".
@@ Appendix F LLM Output Example
{"Literature Review": "CRISPR-Cas systems have emerged as a powerful tool in the field of biosensing, offering unprecedented capabilities for detecting nucleic acids, proteins, and small molecules. This review explores the technical developments and application prospects of CRISPR-based biosensing strategies, highlighting their potential to revolutionize diagnostics and molecular detection.The core biochemical properties of CRISPR-Cas systems provide the foundation for their application in biosensing. These properties include diverse reaction temperatures, programmability in design, high reaction efficiency, and recognition specificity. Chen et al. (2018) demonstrated the versatility of CRISPR-Cas9 in detecting DNA targets across a wide temperature range, from 20°C to 60°C, enabling its use in various environmental conditions. The programmability of CRISPR systems allows for the easy design of guide RNAs (gRNAs) to target specific sequences, as shown by Gootenberg et al. (2017) in their development of the SHERLOCK (Specific High-sensitivity Enzymatic Reporter unLOCKing) platform for nucleic acid detection.The high reaction efficiency of CRISPR-Cas systems contributes to their sensitivity in biosensing applications. Li et al. (2019) reported a CRISPR-Cas12a-based assay capable of detecting attomolar concentrations of DNA targets, demonstrating the potential for ultra-sensitive detection. The recognition specificity of CRISPR systems allows for precise discrimination between closely related sequences, as evidenced by the work of Myhrvold et al. (2018), who developed a CRISPR-based diagnostic platform capable of distinguishing between Zika and Dengue virus strains with single-base resolution.Recent efforts have focused on improving these core properties to enhance the performance of CRISPR-based biosensors. For example, Harrington et al. (2018) engineered Cas12a variants with improved activity and specificity, expanding the range of targetable sequences and reducing off-target effects. Similarly, Gootenberg et al. (2018) developed an enhanced version of the SHERLOCK system, incorporating lateral flow readouts and multiplexing capabilities, further improving its utility in point-of-care diagnostics.One of the key areas of technical development in CRISPR-based biosensing is improving sensitivity and quantification capabilities. Wang et al. (2020) introduced a CRISPR-Cas12a-based digital PCR method that achieved attomolar sensitivity and precise quantification of target nucleic acids. This approach combines the specificity of CRISPR recognition with the quantitative power of digital PCR, opening new avenues for highly sensitive and accurate molecular diagnostics.Multiplex detection is another important aspect of CRISPR biosensing development. Ackerman et al. (2020) demonstrated a CRISPR-Cas13-based multiplexed assay capable of simultaneously detecting and distinguishing multiple SARS-CoV-2 variants, showcasing the potential of CRISPR technology in rapid and accurate pathogen identification. The ability to detect multiple targets in a single reaction not only saves time and resources but also provides more comprehensive diagnostic information.Efforts to develop convenient one-pot assays have led to simplified and more user-friendly CRISPR-based detection methods. Joung et al. (2020) reported a one-pot, CRISPR-Cas12-based assay for SARS-CoV-2 detection that could be performed at a single temperature, eliminating the need for complex thermal cycling equipment. This advancement brings CRISPR-based diagnostics closer to point-of-care applications, where simplicity and ease of use are crucial.The creation of advanced sensors has further expanded the capabilities of CRISPR-based biosensing. Bruch et al. (2019) developed a CRISPR-Cas13a-based electrochemical sensor for nucleic acid detection, demonstrating the potential for integrating CRISPR technology with electronic readout systems. This approach offers the possibility of developing portable, electronic CRISPR-based diagnostic devices for rapid and sensitive detection in resource-limited settings.While CRISPR-based biosensing has shown great promise, there are still obstacles to its widespread commercial application. These include challenges in sample preparation, the need for standardization across different CRISPR systems and assay formats, and regulatory considerations for diagnostic use. Addressing these issues will be crucial for the successful translation of CRISPR biosensing technology from the laboratory to real-world applications.Looking ahead, the future of CRISPR-based biosensing appears bright, with numerous opportunities for further development and application. Potential directions include the integration of CRISPR biosensors with microfluidic devices for automated sample processing and analysis, the development of CRISPR-based wearable sensors for continuous health monitoring, and the application of machine learning algorithms to improve the design and performance of CRISPR biosensing systems.In conclusion, CRISPR-based biosensing strategies represent a rapidly evolving field with immense potential to transform molecular detection and diagnostics. As researchers continue to refine and expand upon the core capabilities of CRISPR systems, we can expect to see increasingly sensitive, specific, and versatile biosensing applications emerge, with far-reaching implications for healthcare, environmental monitoring, and beyond.", "References": [{"title": "CRISPR–Cas12a target binding unleashes indiscriminate single-stranded DNase activity", "authors": "Chen, Janice S., Ma, Enbo, Harrington, Lucas B., Da Costa, Maria, Tian, Xiang, Palefsky, Joel M., Doudna, Jennifer A.", "journal": "Science", "year": "2018", "volumes": "360", "first page": "436", "last page": "439"},{"title": "Nucleic acid detection with CRISPR-Cas13a/C2c2", "authors": "Gootenberg, Jonathan S., Abudayyeh, Omar O., Lee, Jeong Wook, Essletzbichler, Patrick, Dy, Aaron J., Joung, Julia, Verdine, Vanessa, Donghia, Nina, Daringer, Nichole M., Freije, Catherine A.", "journal": "Science", "year": "2017", "volumes": "356", "first page": "438", "last page": "442", "DOI": "10.1126/science.aam9321"}, {"title": "CRISPR-Cas12a-assisted nucleic acid detection", "authors": "Li, Suwei, Cheng, Qingmei, Wang, Jianming, Li, Xiaoyu, Zhang, Zhiwei, Gao, Shan, Cao, Rong, Zhao, Guoping, Wang, Jin", "journal": "Cell Discovery", "year": "2019", "volumes": "5", "first page": "1", "last page": "4"}, {"title": "Field-deployable viral diagnostics using CRISPR-Cas13", "authors": "Myhrvold, Cameron, Freije, Catherine A., Gootenberg, Jonathan S., Abudayyeh, Omar O., Metsky, Hayden C., Durbin, Ann F., Kellner, Max J., Tan, Amanda L., Paul, Lauren M., Parham, Leda A.", "journal": "Science", "year": "2018", "volumes": "360", "first page": "444", "last page": "448"}, {"title": "Enhanced proofreading governs CRISPR–Cas9 targeting accuracy", "authors": "Harrington, Lucas B., Paez-Espino, David, Staahl, Brett T., Chen, Janice S., Ma, Enbo, Kyrpides, Nikos C., Doudna, Jennifer A.", "journal": "Nature", "year": "2018", "volumes": "563", "first page": "621", "last page": "625"}, {"title": "Multiplexed and portable nucleic acid detection platform with Cas13, Cas12a, and Csm6", "authors": "Gootenberg, Jonathan S., Abudayyeh, Omar O., Kellner, Max J., Joung, Julia, Collins, James J., Zhang, Feng", "journal": "Science", "year": "2018", "volumes": "360", "first page": "439", "last page": "444"}, {"title": "Ultrasensitive and visual detection of SARS-CoV-2 using all-in-one dual CRISPR-Cas12a assay", "authors": "Wang, Xiaoxia, Zhong, Minjie, Liu, Yue, Ma, Pengfei, Dang, Lei, Meng, Qing, Wan, Wanying, Ma, Xiaowei, Liu, Jing, Yang, Guohua", "journal": "Nature Communications", "year": "2020", "volumes": "11", "first page": "4711", "last page": "4711"}, {"title": "Detection of SARS-CoV-2 with SHERLOCK One-Pot Testing", "authors": "Joung, Julia, Ladha, Alim, Saito, Makoto, Kim, Nam-Gyun, Woolley, Ann E., Segel, Michael, Barretto, Robert P. J., Ranu, Antonija, Macrae, Rhiannon K., Faure, Guilhem", "journal": "New England Journal of Medicine", "year": "2020", "volumes": "383", "first page": "1492", "last page": "1494"}, {"title": "CRISPR-Cas13-based electrochemical biosensing of viral RNA: Application to detection of SARS-CoV-2", "authors": "Bruch, Richard, Baaske, Johannes, Chatelle, Claire, Meirich, Maren, Madlener, Sibylle, Weber, Wilfried, Dincer, Can, Urban, Gerald A.", "journal": "Angewandte Chemie International Edition", "year": "2019", "volumes": "58", "first page": "17571", "last page": "17575"}, {"title": "Scalable and robust SARS-CoV-2 testing in an academic center", "authors": "Ackerman, Cheri M., Myhrvold, Cameron, Thakku, Shiv G., Freije, Catherine A., Metsky, Hayden C., Yang, David K., Ye, Simon H., Boehm, Chloe K., Kosoko-Thoroddsen, Tinna-Solveig F., Kehe, Jared", "journal": "Nature Biotechnology", "year": "2020", "volumes": "38", "first page": "927", "last page": "931"}]}

### v5

@@ 3.3 Evaluation Metrics
$\text{True}(r_{n}^{g})=\begin{cases}1&\text{if }\left(e_{0}=1\text{ and }\sum_{i=1}^{6}e_{i}\geq 1\right)\\ &\text{or }\left(e_{0}=0\text{ and }\sum_{i=1}^{6}e_{i}\geq 3\right)\\ 0&\text{otherwise}\end{cases}\vskip-5.69054pt$ (1)
@@
$Precision_{(p_{i})}=\frac{1}{N}\sum_{n=0}^{N}\text{True}(r_{n}^{g})\vskip-5.69054pt$ (2)
$Preicison=\frac{1}{M}\sum_{i=0}^{M}Precison{(p_{i})}\vskip-2.84526pt$ (3)
@@
$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}^{(n)}\vskip-5.69054pt$ (4)
[…]
@@
$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}\vskip-5.69054pt$ (7)
@@
$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}\vskip-5.69054pt$ (8)
@@ 4.2 Main Results
Table 1: The experimental results of the five LLMs in Reference Generation. $S_{t}$” refers to the title search rate as defined in Eq 4, while “P” represents the Precision, “R” denotes the Recall. “first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
@@
Table 3: The experimental results of the four LLMs in Review Composition. $S_{t}$” refers to the title retrieval rate as defined in Eq 4. While “P” represents the Precision, “R” denotes the Recall. “KPR” means the Key Point Recall rate.“first author” refers to when evaluating the accuracy of references, the author dimension only comparing the first author.
@@ Appendix A Prompts for Tasks
Prompt 1 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to find the 10 studies that are most relevant to the research topic provided in the ”Title” and the ”Keywords” below.
Please cite the studies according to the following JSON format. There is no need to provide any explanation before or after the JSON output. Ensure that the ”authors” field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as ”et al.”. { ”References”: [ { ”title”: ””, ”authors”: ””, ”journal”: ””, ”year”: ””, ”volumes”: ””, ”first page”: ””, ”last page”: ””, } ] }
@@
Prompt 2 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write an abstract according to the research topic provided in the ”Title” and the ”Keywords” below. Please write the abstract for about xx words, according to the JSON format as follows. There is no need to provide any explanation before or after the JSON output. {”Abstract”: ””}
@@
Prompt 3 Imagine you are an experienced academic researcher with access to a vast library of scientific literature. I would like you to write a literature review according to the research topic provided in the ”Title”, “Abstract” and ”Keywords” below. The literature review should be about 1000 words long. I would like you to back up claims by citing previous studies (with a total of 10 citations in the literature review). The output should be in JSON format as follows: { ”Literature Review”: ”xxx”, ”References”: [ { ”title”: ””, ”authors”: ””, ”journal”: ””, ”year”: ””, ”volumes”: ””, ”first page”: ””, ”last page”: ””, } ] }
The ”Literature Review” field should be about 1000 words. The ”References” field is a list of 10 references, and ensures that the ”authors” field lists the names of all authors and not exceeding 10 authors, and that there are no duplicate author names nor abbreviations such as ”et al.”.
@@ Appendix F LLM Output Example
{”Literature Review”: ”CRISPR-Cas systems have emerged as a powerful tool in the field of biosensing, offering unprecedented capabilities for detecting nucleic acids, proteins, and small molecules. This review explores the technical developments and application prospects of CRISPR-based biosensing strategies, highlighting their potential to revolutionize diagnostics and molecular detection.The core biochemical properties of CRISPR-Cas systems provide the foundation for their application in biosensing. These properties include diverse reaction temperatures, programmability in design, high reaction efficiency, and recognition specificity. Chen et al. (2018) demonstrated the versatility of CRISPR-Cas9 in detecting DNA targets across a wide temperature range, from 20°C to 60°C, enabling its use in various environmental conditions. The programmability of CRISPR systems allows for the easy design of guide RNAs (gRNAs) to target specific sequences, as shown by Gootenberg et al. (2017) in their development of the SHERLOCK (Specific High-sensitivity Enzymatic Reporter unLOCKing) platform for nucleic acid detection.The high reaction efficiency of CRISPR-Cas systems contributes to their sensitivity in biosensing applications. Li et al. (2019) reported a CRISPR-Cas12a-based assay capable of detecting attomolar concentrations of DNA targets, demonstrating the potential for ultra-sensitive detection. The recognition specificity of CRISPR systems allows for precise discrimination between closely related sequences, as evidenced by the work of Myhrvold et al. (2018), who developed a CRISPR-based diagnostic platform capable of distinguishing between Zika and Dengue virus strains with single-base resolution.Recent efforts have focused on improving these core properties to enhance the performance of CRISPR-based biosensors. For example, Harrington et al. (2018) engineered Cas12a variants with improved activity and specificity, expanding the range of targetable sequences and reducing off-target effects. Similarly, Gootenberg et al. (2018) developed an enhanced version of the SHERLOCK system, incorporating lateral flow readouts and multiplexing capabilities, further improving its utility in point-of-care diagnostics.One of the key areas of technical development in CRISPR-based biosensing is improving sensitivity and quantification capabilities. Wang et al. (2020) introduced a CRISPR-Cas12a-based digital PCR method that achieved attomolar sensitivity and precise quantification of target nucleic acids. This approach combines the specificity of CRISPR recognition with the quantitative power of digital PCR, opening new avenues for highly sensitive and accurate molecular diagnostics.Multiplex detection is another important aspect of CRISPR biosensing development. Ackerman et al. (2020) demonstrated a CRISPR-Cas13-based multiplexed assay capable of simultaneously detecting and distinguishing multiple SARS-CoV-2 variants, showcasing the potential of CRISPR technology in rapid and accurate pathogen identification. The ability to detect multiple targets in a single reaction not only saves time and resources but also provides more comprehensive diagnostic information.Efforts to develop convenient one-pot assays have led to simplified and more user-friendly CRISPR-based detection methods. Joung et al. (2020) reported a one-pot, CRISPR-Cas12-based assay for SARS-CoV-2 detection that could be performed at a single temperature, eliminating the need for complex thermal cycling equipment. This advancement brings CRISPR-based diagnostics closer to point-of-care applications, where simplicity and ease of use are crucial.The creation of advanced sensors has further expanded the capabilities of CRISPR-based biosensing. Bruch et al. (2019) developed a CRISPR-Cas13a-based electrochemical sensor for nucleic acid detection, demonstrating the potential for integrating CRISPR technology with electronic readout systems. This approach offers the possibility of developing portable, electronic CRISPR-based diagnostic devices for rapid and sensitive detection in resource-limited settings.While CRISPR-based biosensing has shown great promise, there are still obstacles to its widespread commercial application. These include challenges in sample preparation, the need for standardization across different CRISPR systems and assay formats, and regulatory considerations for diagnostic use. Addressing these issues will be crucial for the successful translation of CRISPR biosensing technology from the laboratory to real-world applications.Looking ahead, the future of CRISPR-based biosensing appears bright, with numerous opportunities for further development and application. Potential directions include the integration of CRISPR biosensors with microfluidic devices for automated sample processing and analysis, the development of CRISPR-based wearable sensors for continuous health monitoring, and the application of machine learning algorithms to improve the design and performance of CRISPR biosensing systems.In conclusion, CRISPR-based biosensing strategies represent a rapidly evolving field with immense potential to transform molecular detection and diagnostics. As researchers continue to refine and expand upon the core capabilities of CRISPR systems, we can expect to see increasingly sensitive, specific, and versatile biosensing applications emerge, with far-reaching implications for healthcare, environmental monitoring, and beyond.”, ”References”: [{”title”: ”CRISPR–Cas12a target binding unleashes indiscriminate single-stranded DNase activity”, ”authors”: ”Chen, Janice S., Ma, Enbo, Harrington, Lucas B., Da Costa, Maria, Tian, Xiang, Palefsky, Joel M., Doudna, Jennifer A.”, ”journal”: ”Science”, ”year”: ”2018”, ”volumes”: ”360”, ”first page”: ”436”, ”last page”: ”439”},{”title”: ”Nucleic acid detection with CRISPR-Cas13a/C2c2”, ”authors”: ”Gootenberg, Jonathan S., Abudayyeh, Omar O., Lee, Jeong Wook, Essletzbichler, Patrick, Dy, Aaron J., Joung, Julia, Verdine, Vanessa, Donghia, Nina, Daringer, Nichole M., Freije, Catherine A.”, ”journal”: ”Science”, ”year”: ”2017”, ”volumes”: ”356”, ”first page”: ”438”, ”last page”: ”442”, ”DOI”: ”10.1126/science.aam9321”}, {”title”: ”CRISPR-Cas12a-assisted nucleic acid detection”, ”authors”: ”Li, Suwei, Cheng, Qingmei, Wang, Jianming, Li, Xiaoyu, Zhang, Zhiwei, Gao, Shan, Cao, Rong, Zhao, Guoping, Wang, Jin”, ”journal”: ”Cell Discovery”, ”year”: ”2019”, ”volumes”: ”5”, ”first page”: ”1”, ”last page”: ”4”}, {”title”: ”Field-deployable viral diagnostics using CRISPR-Cas13”, ”authors”: ”Myhrvold, Cameron, Freije, Catherine A., Gootenberg, Jonathan S., Abudayyeh, Omar O., Metsky, Hayden C., Durbin, Ann F., Kellner, Max J., Tan, Amanda L., Paul, Lauren M., Parham, Leda A.”, ”journal”: ”Science”, ”year”: ”2018”, ”volumes”: ”360”, ”first page”: ”444”, ”last page”: ”448”}, {”title”: ”Enhanced proofreading governs CRISPR–Cas9 targeting accuracy”, ”authors”: ”Harrington, Lucas B., Paez-Espino, David, Staahl, Brett T., Chen, Janice S., Ma, Enbo, Kyrpides, Nikos C., Doudna, Jennifer A.”, ”journal”: ”Nature”, ”year”: ”2018”, ”volumes”: ”563”, ”first page”: ”621”, ”last page”: ”625”}, {”title”: ”Multiplexed and portable nucleic acid detection platform with Cas13, Cas12a, and Csm6”, ”authors”: ”Gootenberg, Jonathan S., Abudayyeh, Omar O., Kellner, Max J., Joung, Julia, Collins, James J., Zhang, Feng”, ”journal”: ”Science”, ”year”: ”2018”, ”volumes”: ”360”, ”first page”: ”439”, ”last page”: ”444”}, {”title”: ”Ultrasensitive and visual detection of SARS-CoV-2 using all-in-one dual CRISPR-Cas12a assay”, ”authors”: ”Wang, Xiaoxia, Zhong, Minjie, Liu, Yue, Ma, Pengfei, Dang, Lei, Meng, Qing, Wan, Wanying, Ma, Xiaowei, Liu, Jing, Yang, Guohua”, ”journal”: ”Nature Communications”, ”year”: ”2020”, ”volumes”: ”11”, ”first page”: ”4711”, ”last page”: ”4711”}, {”title”: ”Detection of SARS-CoV-2 with SHERLOCK One-Pot Testing”, ”authors”: ”Joung, Julia, Ladha, Alim, Saito, Makoto, Kim, Nam-Gyun, Woolley, Ann E., Segel, Michael, Barretto, Robert P. J., Ranu, Antonija, Macrae, Rhiannon K., Faure, Guilhem”, ”journal”: ”New England Journal of Medicine”, ”year”: ”2020”, ”volumes”: ”383”, ”first page”: ”1492”, ”last page”: ”1494”}, {”title”: ”CRISPR-Cas13-based electrochemical biosensing of viral RNA: Application to detection of SARS-CoV-2”, ”authors”: ”Bruch, Richard, Baaske, Johannes, Chatelle, Claire, Meirich, Maren, Madlener, Sibylle, Weber, Wilfried, Dincer, Can, Urban, Gerald A.”, ”journal”: ”Angewandte Chemie International Edition”, ”year”: ”2019”, ”volumes”: ”58”, ”first page”: ”17571”, ”last page”: ”17575”}, {”title”: ”Scalable and robust SARS-CoV-2 testing in an academic center”, ”authors”: ”Ackerman, Cheri M., Myhrvold, Cameron, Thakku, Shiv G., Freije, Catherine A., Metsky, Hayden C., Yang, David K., Ye, Simon H., Boehm, Chloe K., Kosoko-Thoroddsen, Tinna-Solveig F., Kehe, Jared”, ”journal”: ”Nature Biotechnology”, ”year”: ”2020”, ”volumes”: ”38”, ”first page”: ”927”, ”last page”: ”931”}]}